A truck requires 108 litres of diesel for covering a distance of 594 km. How much diesel will be required by the truck to cover a distance of 1650 km? 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Diesel required for 594 km = 108 litres
Diesel required for 1 km = 108 / 594
= 2 / 11 litre
So, diesel required for 1650 km = (2 / 11) × 1650 = 300 litres
∴Diesel required by the truck to cover a distance of 1650 km is 300 litres

Thanks